Talent show tickets are $3 for adults and $2 for students. Ticket sales of at least $600 are needed. At least 5 times as many students as adults attended. Write a system of equations to find the number of adult and student tickets that needed to be sold. Let x represent the number of adult tickets and y represent the number of student tickets. How many of each type of ticket needs to be sold?
600 = 3x + 2y
5x+x= y
6x= y

600 = 3x + 2( 6x)
600 = 15x
40 = x
6x= y
6(40)= y
240= y

600 = 3x + 2y
600 = 3(40)+ 2(240)
600 = 120 + 480
600 = 600

x=40
y=240

Answer by silencehurts(5)

600 = 3x + 2y
5x+x= y
6x= y

600 = 3x + 2( 6x)
600 = 15x
40 = x
6x= y
6(40)= y
240= y

600 = 3x + 2y
600 = 3(40)+ 2(240)
600 = 120 + 480
600 = 600

x=40
y=240